# Physics of Strong Magnetism with eXTP

### v1

@@ Abstract
In this paper we present the science potential of the enhanced X-ray Timing and Polarimetry (eXTP) mission, in its new configuration, for studies of strongly magnetized compact objects. We discuss the scientific potential of eXTP for QED studies, especially leveraging on the recent observations made with the NASA IXPE mission. Given eXTP’s unique combination of timing, spectroscopy, and polarimetry, we focus on the perspectives for physics and astrophysics studies of strongly magnetized compact objects, such as magnetars and accreting X-ray pulsars. Developed by an international Consortium led by the Institute of High Energy Physics of the Chinese Academy of Sciences, the eXTP mission is expected to launch in early 2030.
@@
Neutron Stars, QED, Magnetars, Accreting Pulsars, eXTP1 Introduction
@@
The enhanced X-ray Timing and Polarimetry mission (eXTP) is a science mission designed to study the state of matter under extreme conditions of density, gravity, and magnetism [242, 243]. In the new baseline, eXTP will be launched in early 2030. The payload of the mission consists of three main instruments: the Spectroscopic Focusing Array (SFA), the Polarimetry Focusing Array (PFA) and the Wide-band and Wide-field Camera (W2C). Here we provide a brief introduction to the scientific payloads and refer to xxx need add ref. for a detailed description.
@@
The SFA consists of five SFA-T (where T denotes Timing) X-ray focusing telescopes covering the energy range –, featuring a total effective area of at and at . The designed angular resolution of the SFA is (HPD) with a field of view (FoV). The SFA-T are equipped with silicon-drift detectors (SDDs), which combine good spectral resolution ( 180 eV at 1.5 keV) with very short dead time and a high time resolution of . They are therefore well-suited for studies of compact objects emitting X-rays at the shortest time scales [243]. The SFA array also includes one unit of the SFA-I (where I signifies Imaging) telescope equipped with pn-CCD detectors (p-n junction charged coupled device), to enhance imaging capabilities, which would supply strong upper limits to detect weak and extended sources. The expected FoV of SFA-I is . Therefore, the overall sensitivity of SFA could reach around for an exposure time of . Since it is not excluded that the SFA might in the end include six SFA-T units, simulations presented here have taken this possibility into consideration.
@@
The PFA features three identical telescopes, with an angular resolution better than (HPD) in a FoV, and a total effective area of at (considering the detector efficiency). Polarization measurements are carried out by gas pixel detectors (GPDs) working at 2—- 8 keV with an expected energy resolution of 20% at 6 keV and a time resolution better than [29, 14, 15, 16, 243]. The instrument reaches an expected minimum detectable polarization (MDP) at confidence level () of about in for a milliCrab-like source.
@@
The W2C is a secondary instrument of the science payload, featuring a coded mask camera with a FoV of approximately 1500 square degrees (Full-Width Zero Response, FWZR). The instrument achieves a sensitivity of (1 s exposure) across the 10–600 keV energy range, with an angular resolution of and an energy resolution better than at 60 keV.
@@
It is fascinating to study the behavior of matter and radiation in strong magnetic fields, especially at strengths near or even far beyond the quantum critical field . Magnetars (observationally identified as anomalous X-ray pulsars, AXPs, and soft-gamma repeaters, SGRs) and pulsars in high-mass X-ray binaries supply astrophysical laboratories to reveal the nature of strongly magnetized sources. In the previous eXTP white paper dedicated to strong magnetism science, many topics have been discussed, such as the potential of the mission to investigate quantum electrodynamic (QED) effects, or the bursting activity and accretion mechanisms in neutron stars [NSs, 174]. However, since that publication, the design of the eXTP mission has evolved. Moreover, from the scientific point of view the situation underwent a rapid change, due to the operation of Polar-light and IXPE which achieved polarization measurements in X-rays, and collected data that lead to a quantum leap in our understanding of strongly magnetized sources [51, 241, 217, 73, 194, 150, 169, 181].
@@ 2.1 QED
where is the star magnetic field strength at the pole, is the NS radius, and is the photon energy. An important point is that, independently of the complexity of the magnetic field topology near the star surface, at such a large distance from the star surface the magnetic field can be reasonably approximated as a pure dipole (any multipolar component would have substantially decayed). Therefore, escaping radiation crosses at different positions, where, however, the orientation of the magnetic field is roughly the same. As a result, the polarization state of the observed radiation is expected to closely follow the O/X-mode ratio at the emission, allowing one to gain insight into the emission mechanism that produced the signal. In contrast, if QED effects were not present around the star, the direction of the polarization vectors would not change as radiation travels to the observer, while the magnetic field does. This implies that the normal polarization modes would be washed out, thus substantially reducing the observed polarization degree (PD), even for photons originally emitted as polarized at [196, 63].
@@
Among NSs, magnetars, with magnetic fields in the range – [i.e. – times stronger than those of the most common rotation-powered pulsars, 142] and emission peaked in the X-rays [216], are the best targets for observing this QED effect at work. In these sources, is about (as shown by equation 2), so that the expected polarization pattern closely follows that at the emission. The same considerations hold also for other classes of strongly magnetized NSs, like the X-ray dim isolated neutron stars (for which ).
@@
Nevertheless, the difference expected in the polarization signal when computed with and without taking into account vacuum birefringence tends to reduce as the emitting region shrinks. In fact, if photons originate from a localized spot on the star surface, the mutual orientation of their polarization vectors with respect to the star’s magnetic field is basically the same even at the emission point (because the magnetic field is almost uniform across the spot), and the polarization modes are not changing much along the photon path. Hence, in this case, the intrinsic polarization degree (where for intrinsic we mean the polarization degree at the emission point, so at the star surface) should not be much different from the observed one, regardless of the presence of QED effects [251]. This is shown in Fig. 2, where the observed degree of polarization is plotted as a function of the size of the emitting region, for a computation made either with (QED-on) or without (QED-off) accounting for QED effects (in this example, radiation is assumed to be intrinsically polarized at about in the X mode). So, for the detection of a high polarization degree to be a proof of vacuum birefringence, observations should target highly magnetized NSs for which the X-ray radiation originates from a sufficiently large emission region (typically of the star surface), which implies a low pulsed fraction.
@@
Transient magnetars (hereafter TMs), if observed relatively soon after the onset of an outburst event, satisfy this requirement (see Section 2.4.2 for further details); but also some of the known persistent magnetars do. Among the magnetars observed so far by IXPE, the only one exhibiting a quite low pulsed fraction, , is SGR 180620 [217]. However, despite the long exposure time (), the low flux level, in the – range prevented a secure measurement of the polarization properties with IXPE. In this respect, eXTP, with its increased effective area, promises a much better sensitivity, ensuring a secure detection with similar or shorter exposure times. An example is shown in Fig. 3, where we report the expected degree of polarization for a simulated observation of SGR 180620 made with eXTP, compared to an IXPE observation of the same duration. The calculation is based on the same model discussed in [217], i.e. thermal emission from a condensed surface region located along the magnetic equator, reprocessed by resonant Compton scattering (RCS) in the inner magnetosphere (see §2.2.2). The plots clearly show that an observation performed by eXTP (central panel) will lead to a significant improvement of the IXPE measurement (left panel), allowing for a signal detection above in all the energy bins across the entire – band, for the model self-consistently computed considering QED effects. This will provide a confirmation of the theoretical interpretation of the emission, which could not be firmly obtained with IXPE because the polarization was measured only in a single energy bin at confidence level. In order to show how eXTP polarization measurements may be used to test vacuum birefringence, we performed an analogous simulation without including QED effects. As the right panel of Fig. 3 shows, the polarization in this case is expected to be much lower than in the previous case in all energy bins.
@@
An alternative way to probe whether vacuum birefringence is indeed at work around magnetars is by looking at how the polarization properties of the observed X-ray emission change with the star’s rotational phase. IXPE observations have shown that, for persistent magnetars, the polarization degree and angle exhibit different phase-dependent behaviors [198, 241, 73, 150], with the former following the light curve and the latter oscillating according to the rotating vector model [RVM; 165, 160]. The behavior of the polarization degree can be explained if the polarization signal originates close to the star surface, and in such a case it will give insight into the emission geometry. Instead, we investigate the behavior of the polarization angle, which is completely unexpected for persistent magnetars, given that the RVM assumes dipolar magnetic fields, like those expected in rotation-powered pulsars. In magnetars, instead, the magnetic field close to the surface is likely far from being dipolar. The only way out of this conundrum is to invoke vacuum birefringence. If the polarization-limiting radius is large enough (as in magnetars, see equation 2), the pattern observed at infinity in the polarization degree reflects that of the emission near the surface, while the observed polarization vector mimics the megnetic field topology at , where only the dipolar component is important.
The capability of eXTP to make phase-resolved observations of persistent magnetars with a large accuracy, and to expand the number of observable targets including other classes of sources, is crucial to reaffirm and consolidate this conclusion, which is currently based on few preliminary results achieved up to now with IXPE.
@@ 2.2.1.1 Proton CRSFs in magnetar spectra
Some of the currently known magnetar sources, such as SGR 04185729 and AXP 1E 2259586, exhibit absorption features at X-ray energies, with peculiar variability across the rotational phase. SGR 04185729 entered an outburst phase in 2009, when the X-ray flux increased to , and at that time an absorption line was detected in the spectrum by XMM-Newton, RXTE, and Swift, with a line energy in the – range [201]. An absorption line was also visible in the spectrum of AXP 1E 2259586, in two distinct sets of XMM-Newton observations, performed in 2002 and 2014, centered at an energy of [154]. The presence of this spectral line has been further confirmed by a joint XMM-Newton and IXPE observation performed in 2023 [73]. Furthermore, spectral lines with similar properties have also been observed in the AXP Swift J1822.31606 [170], as well as in some X-ray dim isolated neutron stars [XDINSs, 21, 22].
@@
(indicative of a polarization mode switch) occurs. Heyl
et al. [73] found that these results are compatible with the reprocessing of photons via RCS in a localized magnetic loop. However, due to the low flux of the source, the IXPE observation was not sufficient to constrain the observed quantities to more than , even with an exposure time . The strong increase in the effective area ensured by eXTP, as well as the possibility of performing spectral and polarization observations at the same time from the same source, promise to increase the sensitivity for targets such as AXP 1E 2259586. Such a systematic spectro-polarimetric study of CRSFs in magnetars (and, possibly, in other classes of highly magnetized NSs, such as the XDINSs) will help to identify the nature of the lines, potentially allowing one to map the surface magnetic field strength and topology in these sources. Transient magnetars, during outburst phases, may also exhibit similar lines in their spectra, In this case, their study will allow one to resolve the untwisting of the magnetic loops above the heated surface region during the outburst decay, providing an independent test of the magnetar scenario (see section 2.4.2).
@@ 2.2.1.2 Electron CRSFs in X-ray pulsars spectra
For X-ray pulsars (XRPs), CRSFs due to resonant scattering onto electrons are considered the most direct method for measuring the magnetic field strength in the emission region, that is, in the vicinity of the NS surface. Currently, approximately 40 XRPs are known to exhibit absorption features in their spectra [180, 135]. However, the vast majority of known CRSFs are located in the hard X-ray band, above 20 keV. This bias arises in part due to the complex nature of XRP spectra below 10 keV, where multiple spectral components, including the soft excess, iron emission line, iron edge, and hard continuum, overlap, leading to strong degeneracies that complicate the identification of additional absorption features. As a result, the measured distribution of magnetic field strengths in XRPs is skewed toward higher values. Another challenge in CRSF detection is their potential phase dependence. In some cases, CRSFs appear only within a narrow range of spin phases, making them difficult to detect in phase-averaged spectra. However, several CRSFs below 10 keV have been discovered [39, 132, 133], some of which are phase-transient, a finding that highlights the need for detailed phase-resolved spectral analysis. Conducting such an analysis requires exceptionally high photon statistics, which will be provided by eXTP/SFA. A key advantage of eXTP is its ability to perform simultaneous energy-resolved polarimetry with the PFA instrument in the same energy band. This unique combination will enable an unambiguous identification of CRSFs in soft X-rays by detecting the characteristic 90 degree swing in the polarization position angle, expected due to mode conversion at the electron-cyclotron resonance [138, 179]. This polarization signature will break degeneracies caused by complex spectral continua, providing a robust confirmation of CRSFs and offering new insights into the structure of magnetized atmospheres and accretion columns.

### v2

@@ Abstract
In this paper we present the science potential of the enhanced X-ray Timing and Polarimetry (eXTP) mission, in its new configuration, for studies of strongly magnetized compact objects. We discuss the scientific potential of eXTP for quantum electrodynamic (QED) studies, especially leveraging on the recent observations made with the NASA IXPE mission. Given eXTP’s unique combination of timing, spectroscopy, and polarimetry, we focus on the perspectives for physics and astrophysics studies of strongly magnetized compact objects, such as magnetars and accreting X-ray pulsars. Developed by an international Consortium led by the Institute of High Energy Physics of the Chinese Academy of Sciences, the eXTP mission is expected to launch in early 2030.
@@
Neutron Stars, QED, Magnetars, Accreting Pulsars, eXTP97.60.Jd, 97.80.Jp, 12.20.Fv
1 Introduction
@@
The enhanced X-ray Timing and Polarimetry mission (eXTP) is a science mission designed to study the state of matter under extreme conditions of density, gravity, and magnetism [1, 2]. In the new baseline, eXTP will be launched in early 2030. The payload of the mission consists of three main instruments: the Spectroscopic Focusing Array (SFA), the Polarimetry Focusing Array (PFA) and the Wide-band and Wide-field Camera (W2C). Here we provide a brief introduction to the scientific payloads. For a detailed description and comparison with other missions, please refer to the white paper of the whole eXTP project [3].
@@
The SFA consists of five SFA-T (where T denotes Timing) X-ray focusing telescopes covering the energy range –, featuring a total effective area of at and at . The designed angular resolution of the SFA is (HPD) with a field of view (FoV). The SFA-T are equipped with silicon-drift detectors (SDDs), which combine good spectral resolution ( 180 eV at 1.5 keV) with very short dead time and a high time resolution of . They are therefore well-suited for studies of compact objects emitting X-rays at the shortest time scales and even for the bright X-ray sources up to [2]. The SFA array also includes one unit of the SFA-I (where I signifies Imaging) telescope equipped with pn-CCD detectors (p-n junction charged coupled device), to enhance imaging capabilities, which would supply strong upper limits to detect weak and extended sources, though suffering the pile-up effects when observing bright sources. The expected FoV of SFA-I is . Therefore, the overall sensitivity of SFA could reach around for an exposure time of 1 million seconds (). Since it is not excluded that the SFA might in the end include six SFA-T units, simulations presented here have taken this possibility into consideration.
@@
The PFA features three identical telescopes, with an angular resolution better than (HPD) in a FoV, and a total effective area of at (considering the detector efficiency). Polarization measurements are carried out by gas pixel detectors (GPDs) working at 2–8 keV with an expected energy resolution of 20% at 6 keV and a time resolution better than [4, 5, 6, 7, 2]. The instrument reaches an expected minimum detectable polarization (MDP) at confidence level () of about in for a milliCrab-like source as shown in Figure 1. In the current MDP calculation, the background is neglected under the assumption that the source is significantly brighter than the background. When the background cannot be neglected for weak sources (0.01 mCrab, ), the MDP changes from 20% to 29% with an exposure of 1 Ms, which is inadequate for confident measurements.
@@
The W2C is a secondary instrument of the science payload, featuring a coded mask camera with a FoV of approximately 1500 square degrees (Full-Width Zero Response, FWZR). The instrument achieves a sensitivity of (1 s exposure) across the 30–600 keV energy range, with an angular resolution of and an energy resolution better than at 60 keV.
@@
It is fascinating to study the behavior of matter and radiation in strong magnetic fields, especially at strengths near or even far beyond the quantum critical field . Magnetars (observationally identified as anomalous X-ray pulsars, AXPs, and soft-gamma repeaters, SGRs) and pulsars in high-mass X-ray binaries supply astrophysical laboratories to reveal the nature of strongly magnetized sources. In the previous eXTP white paper dedicated to strong magnetism science, many topics have been discussed, such as the potential of the mission to investigate quantum electrodynamic (QED) effects, or the bursting activity and accretion mechanisms in neutron stars [NSs, 8]. However, since that publication, the design of the eXTP mission has evolved. Moreover, from the scientific point of view the situation underwent a rapid change, due to the operation of Polar-light and IXPE which achieved polarization measurements in X-rays, and collected data that lead to a quantum leap in our understanding of strongly magnetized sources [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@ 2.1 QED
where is the star magnetic field strength at the pole, is the NS radius, and is the photon energy. An important point is that, independently of the complexity of the magnetic field topology near the star surface, according to the most commonly-accepted scenario the magnetic field can be reasonably approximated as a pure dipole at such a large distance from the star surface (any multipolar, component would have substantially decayed). Therefore, escaping radiation crosses at different positions, where, however, the orientation of the magnetic field is roughly the same. As a result, the polarization state of the observed radiation is expected to closely follow the O/X-mode ratio at the emission, allowing one to gain insight into the emission mechanism that produced the signal. In contrast, if QED effects were not present around the star, the direction of the polarization vectors would not change as radiation travels to the observer, while the magnetic field does. This implies that the normal polarization modes would be washed out, thus substantially reducing the observed polarization degree (PD), even for photons originally emitted as polarized at [27, 28].
If, on the other hand, the magnetic field at the surface is dominated by higher-older multipoles, the polarization-limiting radius may assume values different from those provided by equation 2. By way of example, Fig. 2 shows the behavior of as a function of the (total) surface magnetic field and the photon energy for a pure dipole, dipole quadrupole times as strong at the surface, and dipole octupole times as strong at the surface. Generally, the polarization-limiting radius tends to decrease when the field is dominated by higher-order multipoles. Nevertheless, at typical magnetar field strengths () and for photon energies within the eXTP PDA working band (–), attains quite large values ( in the quadrupolar configuration and in the octupolar one). Despite the presence of multipolar components changes , at such large distances from the star the dipole is anyway dominant. In this situation, a substantial difference in the polarization degree with or without considering QED effects (as described above) is still expected. However, clear effects on the polarization angle behavior, as a result of the more complex magnetic field topology, should be visible (as discussed in detail in the following).
@@
Among NSs, magnetars, with magnetic fields in the range – [i.e. – times stronger than those of the most common rotation-powered pulsars, 29] and emission peaked in the X-rays [30], are the best targets for observing this QED effect at work. In these sources, is about (as shown by equation 2), so that the expected polarization pattern closely follows that at the emission. The same considerations hold also for other classes of strongly magnetized NSs, like the X-ray dim isolated neutron stars (for which ).
@@
Nevertheless, the difference expected in the polarization signal when computed with and without taking into account vacuum birefringence tends to reduce as the emitting region shrinks. In fact, if photons originate from a very small, localized spot on the star surface, the mutual orientation of their polarization vectors with respect to the star’s magnetic field is basically the same even at the emission point (because the magnetic field is almost uniform across the spot), and the polarization modes are not changing much along the photon path. Hence, in this case, the intrinsic polarization degree (where for intrinsic we mean the polarization degree at the emission point, so at the star surface) should not be much different from the observed one, regardless of the presence of QED effects [32]. This is shown in Fig. 3, where the observed degree of polarization is plotted as a function of the size of the emitting region, for a computation made either with (QED-on) or without (QED-off) accounting for QED effects (in this example, radiation is assumed to be intrinsically polarized at about in the X mode). So, for the detection of a high polarization degree to be a proof of vacuum birefringence, observations should target highly magnetized NSs for which the X-ray radiation originates from a sufficiently large emission region (typically of the star surface), which implies a low pulsed fraction.
@@
Transient magnetars (hereafter TMs), if observed relatively soon after the onset of an outburst event, satisfy this requirement (see Section 2.4.2 for further details); but also some of the known persistent magnetars do. Among the magnetars observed so far by IXPE, the only one exhibiting a quite low pulsed fraction, , is SGR 180620 [11]. However, despite the long exposure time (), the low flux level, in the – range prevented a secure measurement of the polarization properties with IXPE. In this respect, eXTP, with its increased effective area, promises a much better sensitivity, ensuring a secure detection with similar or shorter exposure times. An example is shown in Fig. 4, where we report the expected degree of polarization for a simulated observation of SGR 180620 made with eXTP, compared to an IXPE observation of the same duration. The calculation is based on the same model discussed in [11], i.e. thermal emission from a condensed surface region located along the magnetic equator, reprocessed by resonant Compton scattering (RCS) in the inner magnetosphere (see §2.2.2). The plots clearly show that an observation performed by eXTP (central panel) will lead to a significant improvement of the IXPE measurement (left panel), allowing for a signal detection above in all the energy bins across the entire – band, for the model self-consistently computed considering QED effects ( passing from a level of significance to for a exposure time). This will provide a confirmation of the theoretical interpretation of the emission, which could not be firmly obtained with IXPE because the polarization was measured only in a single energy bin at confidence level. In order to show how eXTP polarization measurements may be used to test vacuum birefringence, we performed an analogous simulation without including QED effects. As the right panel of Fig. 4 shows, the polarization in this case is expected to be much lower than in the previous case in all energy bins.
To better illustrate the improvement granted by eXTP on magnetar polarization measurements, Fig. 5 shows the energy-dependent polarization degree as measured by IXPE for the AXP 1RXS J170849.0400910 [see 10] compared with a simulated observation by eXTP obtained for the same observing time () starting from a theoretical model which gives a similar behavior of the polarization properties as a function of the energy111The model, discussed in Zane et al. [10], assumes the neutron star endowed with a magnetic field with polar strength , with emission from a polar cap covered by a magnetized atmosphere (with tempereature and semiaperture ) and from an equatorial belt with the condensed surface exposed (with temperature , colatitudinal aperture and covering in azimuth).. In the latter case, a remarkable reduction in and errors is evident, confirming that eXTP will enable significantly improved results with respect to IXPE for the same observation time or, alternatively, ensure no loss of significance even with reduced exposure times.
@@
An alternative way to probe whether vacuum birefringence is indeed at work around magnetars is by looking at how the polarization properties of the observed X-ray emission change with the star’s rotational phase. IXPE observations have shown that, for persistent magnetars, the polarization degree and angle exhibit different phase-dependent behaviors [34, 10, 12, 14], with the former following the light curve and the latter oscillating according to the rotating vector model [RVM; 35, 36]. The behavior of the polarization degree can be explained if the polarization signal originates close to the star surface, and in such a case it will give insight into the emission geometry. Instead, we investigate the behavior of the polarization angle, which is completely unexpected for persistent magnetars, given that the RVM assumes dipolar magnetic fields, like those expected in rotation-powered pulsars. In magnetars, instead, the magnetic field close to the surface is likely far from being dipolar. The only way out of this conundrum is to invoke vacuum birefringence. If the polarization-limiting radius is large enough (as in magnetars, see equation 2), the pattern observed at infinity in the polarization degree reflects that of the emission near the surface, while the observed polarization vector mimics the magnetic field topology at , where only the dipolar component is important. The capability of eXTP to make phase-resolved observations of persistent magnetars with a large accuracy, and to expand the number of observable targets including other classes of sources, is crucial to reaffirm and consolidate this conclusion, which is currently based on few preliminary results achieved up to now with IXPE.
The situation would be different assuming more complex magnetic field topologies, with stronger multipolar components superimposed on the (weaker) dipole field (as in the examples provided in Fig. 2). In this cases, however, a substantial deviation of the phase-dependent polarization angle from the RVM is expected, as the axial-symmetry of the dipolar configuration would be definitively broken222 This holds true as long as the dipole and multipolar components do not share the same magnetic axis, which is, however, a quite unlikely situation for young sources such as magnetars [see e.g. 37].. The same can be expected for alternative scenarios assuming, for example, nonlinear expansions of quantum-electrodynamics (like, e.g., within the Born-Infeld scheme discussed in [38]). In this respect, the better sensitivity of the eXTP detectors with respect to previous instrumentation will be certainly capable of resolving such deviations in future observations, allowing for a more thorough understanding of the magnetic field topology of magnetar sources.
@@ 2.2 Spectral appearance
Building upon the comprehensive discussion in the preceding section, which thoroughly addressed the main tests of QED effects attainable with X-ray polarimetry, the ensuing section will delve into how the forthcoming eXTP measurements can help in characterizing the intrinsic physical properties of highly magnetized cosmic sources like magnetars and X-ray pulsars.
@@ 2.2.1.1 Proton CRSFs in magnetar spectra
Some of the currently known magnetar sources, such as SGR 04185729 and AXP 1E 2259586, exhibit absorption features at X-ray energies, with peculiar variability across the rotational phase. SGR 04185729 entered an outburst phase in 2009, when the X-ray flux increased to , and at that time an absorption line was detected in the spectrum by XMM-Newton, RXTE, and Swift, with a line energy in the – range [39]. An absorption line was also visible in the spectrum of AXP 1E 2259586, in two distinct sets of XMM-Newton observations, performed in 2002 and 2014, centered at an energy of [40]. The presence of this spectral line has been further confirmed by a joint XMM-Newton and IXPE observation performed in 2023 [12]. Furthermore, spectral lines with similar properties have also been observed in the AXP Swift J1822.31606 [41], as well as in some X-ray dim isolated neutron stars [XDINSs, 42, 43].
@@
(indicative of a polarization mode switch) occurs. Heyl et al. [12] found that these results are compatible with the reprocessing of photons via RCS in a localized magnetic loop. However, due to the low flux of the source, the IXPE observation was not sufficient to constrain the observed quantities to more than , even with an exposure time . The strong increase in the effective area ensured by eXTP, as well as the possibility of performing spectral and polarization observations at the same time from the same source, promise to increase the sensitivity for targets such as AXP 1E 2259586. Such a systematic spectro-polarimetric study of CRSFs in magnetars (and, possibly, in other classes of highly magnetized NSs, such as the XDINSs) will help to identify the nature of the lines, potentially allowing one to map the surface magnetic field strength and topology in these sources. Transient magnetars, during outburst phases, may also exhibit similar lines in their spectra, In this case, their study will allow one to resolve the untwisting of the magnetic loops above the heated surface region during the outburst decay, providing an independent test of the magnetar scenario (see section 2.4.2).
@@ 2.2.1.2 Electron CRSFs in X-ray pulsars spectra
For X-ray pulsars (XRPs), CRSFs due to resonant scattering onto electrons are considered the most direct method for measuring the magnetic field strength in the emission region, that is, in the vicinity of the NS surface. Currently, approximately 40 XRPs are known to exhibit absorption features in their spectra [44, 45]. However, the vast majority of known CRSFs are located in the hard X-ray band, above 20 keV. This bias arises in part due to the complex nature of XRP spectra below 10 keV, where multiple spectral components, including the soft excess, iron emission line, iron edge, and hard continuum, overlap, leading to strong degeneracies that complicate the identification of additional absorption features. As a result, the measured distribution of magnetic field strengths in XRPs is skewed toward higher values. Another challenge in CRSF detection is their potential phase dependence. In some cases, CRSFs appear only within a narrow range of spin phases, making them difficult to detect in phase-averaged spectra. However, several CRSFs below 10 keV have been discovered [46, 47, 48], some of which are phase-transient, a finding that highlights the need for detailed phase-resolved spectral analysis. Conducting such an analysis requires exceptionally high photon statistics, which will be provided by eXTP/SFA. A key advantage of eXTP is its ability to perform simultaneous energy-resolved polarimetry with the PFA instrument in the same energy band. This unique combination will enable an unambiguous identification of CRSFs in soft X-rays by detecting the characteristic 90 degree swing in the polarization position angle, expected due to mode conversion at the electron-cyclotron resonance [49, 50]. This polarization signature will break degeneracies caused by complex spectral continua, providing a robust confirmation of CRSFs and offering new insights into the structure of magnetized atmospheres and accretion columns. Futhermore, CRSF energies provide a direct probe of the local magnetic field strength, which is sensitive to both the height and latitude of the accretion region and hence crucial for constraining the surface magnetic topology of neutron stars. While the correlation between higher-energy CRSFs ( 10 keV) and accretion rate, driven largely by changes in the accretion column height, is well-established in accreting pulsars [44, 45], the behavior of low-energy ( 10 keV) CRSFs remains unexplored. As CRSF energies trace emission height [51] and phase-resolved polarization measurements inform on the pulsar geometry [52], simultaneous spectroscopy and polarimetry are therefore key to capturing how magnetic field configurations and accretion geometries evolve with accretion rate and time. The unprecedented sensitivity and photon statistics of instruments like SFA and PFA will uniquely enable the first measurements of low-energy CRSF evolution and polarization in accreting pulsars, magnetars and PULXs. Such observations hold the potential to fundamentally advance our understanding of the physics of extreme magnetic fields and accretion processes in neutron stars.